# NUBO: A Transparent Python Package for Bayesian Optimization

### v2

@@ 4 Case study
We present the general workflow for optimizing an expensive-to-evaluate black-box function with NUBO by providing a detailed case study in which a test function with six input dimensions is optimized. This case study demonstrates how the user can specify the parameter input space, generate initial training data, and define and run the Bayesian optimization loop.
@@
    $>$ import torch    $>$    $>$    $>$ torch.manual_seed(123)    $>$ torch.set_printoptions(precision=4, sci_mode=False)A typical objective function optimized with Bayesian optimization is expensive to evaluate and thus not feasible to use in a case study that aims to illustrate how NUBO can be applied. Hence, we will use one of the synthetic test functions provided by NUBO as a surrogate expensive-to-evaluate black-box function. We use the six-dimensional Hartmann function that possesses multiple local and one global minimum. Its input space is bounded by the hyper-rectangle $[0,1]^{6}$. Observational noise, such as measurement error, is represented by adding a small amount of random Gaussian noise to the function output by setting noise_std=0.1. minimise is set to False to transform the minimization into a maximization problem as required for Bayesian optimization with NUBO.
@@
    $>$ from nubo.test_functions import Hartmann6D    $>$    $>$    $>$ black_box = Hartmann6D(noise_std=0.1, minimise=False)With our objective function specified, we can focus on defining the input space. Our objective function has six inputs, all bounded by $[0,1]$. As introduced in Section 3.2, the bounds are defined as a $2\times d$ torch.tensor, where the first row specifies the lower bounds and the second row specifies the upper bounds. This case study also highlights the mixed parameter optimization capabilities of NUBO (see Section 3.2) by assuming that the first input is a discrete parameter restricted to 0.2, 0.4, 0.6, and 0.8. We can implement this by specifying a dictionary, where the key is the input dimension and the value is a list of possible values the input can take, that is {0: [0.0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0]}. Note that indexing starts at zero in Python.
@@
    >>> dims = 6    >>> bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],    ...                        [1., 1., 1., 1., 1., 1.]])    >>> discrete = {0: [0.0, 0.1, 0.2, 0.3, 0.4, 0.5,    ...                 0.6, 0.7, 0.8, 0.9, 1.0]}The Bayesian optimization loop requires initial training data. This is important to train the Gaussian process that tries to emulate the objective function. This case study uses the gen_inputs function introduced in Section 3.3 to generate 30 initial data points from a Latin hypercube design. We round the first input dimension to fit the discrete values specified above as Latin hypercube designs return continuous values. These points are evaluated by the objective function to complete our training data pairs consisting of input parameters x_train and observations y_train.
@@
    $>$ from nubo.utils import gen_inputs    $>$    $>$    $>$ x_train = gen_inputs(num_points=dims * 5,    ...                      num_dims=dims,    ...                      bounds=bounds)    $>$ x_train[:, 0] = torch.round(x_train[:, 0], decimals=1)    $>$ y_train = black_box(x_train)Next, we specify the Bayesian optimization algorithm we plan to use in our optimization loop. We define the bo function that takes our training pairs (x_train, y_train) and returns the next candidate point x_new which is evaluated by the objective function in four steps. First, we set up our surrogate model as the Gaussian process provided by NUBO with a Gaussian likelihood as discussed in Section 3.1. Second, we train the Gaussian process gp with our training data by maximizing the likelihood with the Adam algorithm [2] via the fit_gp function. Here, we set a custom learning rate lr and the number of optimization steps steps. Third, we define an acquisition function that will guide our optimization. Assuming that our objective function allows parallel function evaluations, we compute multi-point batches at each iteration and choose a Monte Carlo acquisition function, in this case MCUpperConfidenceBound. The acquisition function acq is instantiated by providing it with the fitted Gaussian process gp, a value for the trade-off hyper-parameter beta, and the number of Monte Carlo samples used to approximate the acquisition function. For further details, refer to Section 2.2. Fourth, we maximize the acquisition function acq with the multi_sequential function that uses the sequential strategy for computing multiple candidate points. We compute four candidate points at each iteration by setting batch_size=4 and providing the previously specified bounds and discrete values. The Adam optimizer is used as Monte Carlo acquisition functions require a stochastic optimizer due to their inherent randomness introduced by drawing the Monte Carlo samples. The optimizer is initialized at two different initial points chosen as the two points with the highest acquisition value out of 100 potential points sampled from a Latin hypercube design. We chose two initializations to keep the computational overhead within the replication script low. In practice, a higher number of initializations might be beneficial.
@@
    $>$ from nubo.acquisition import MCUpperConfidenceBound    $>$ from nubo.models import GaussianProcess, fit_gp    $>$ from nubo.optimization import multi_sequential    $>$ from gpytorch.likelihoods import GaussianLikelihood    $>$    $>$    $>$ def bo(x_train, y_train):    $>$    $>$     likelihood = GaussianLikelihood()    $>$     gp = GaussianProcess(x_train, y_train, likelihood=likelihood)    $>$    $>$     fit_gp(x_train, y_train, gp=gp, likelihood=likelihood,    ...            lr=0.1, steps=200)    $>$    $>$     acq = MCUpperConfidenceBound(gp=gp, beta=4, samples=128)    $>$    $>$     x_new, _ = multi_sequential(func=acq,    ...                                 method="Adam",    ...                                 batch_size=4,    ...                                 bounds=bounds,    ...                                 discrete=discrete,    ...                                 lr=0.1,    ...                                 steps=200,    ...                                 num_starts=2,    ...                                 num_samples=100)    $>$    $>$     return x_newFinally, we specify the entire optimization loop, that is, a simple for-loop that computes the next batch of candidate points using the defined Bayesian optimization algorithm bo, evaluates the candidate points by the objective function black_box, and adds the new data pairs (x_new, y_new) to the training data. We let the optimization loop run for ten iterations and print all evaluations, where the first six columns are the inputs and the final column is the output from the objective function. The first 30 rows give the initial training data generated by the Latin hypercube design, while the last 40 rows were chosen by the Bayesian optimization algorithm. The results show that NUBO improves upon the initial space-filling design and produces points consistent with the bounds and discrete values that specify the parameter input space.
        $>$ iters = 10        $>$        $>$ for iter in range(iters):        $>$        $>$     x_new = bo(x_train, y_train)        $>$        $>$     y_new = black_box(x_new)        $>$        $>$     x_train = torch.vstack((x_train, x_new))        $>$     y_train = torch.hstack((y_train, y_new))        $>$        $>$ print(torch.hstack([x_train, y_train.reshape(-1, 1)]))
@@
        tensor([[0.2000, 0.6523, 0.1574, 0.7822, 0.3039, 0.8603,  0.1251],                [0.5000, 0.9127, 0.8746, 0.4787, 0.6523, 0.1249,  2.2907],                [0.4000, 0.5638, 0.0459, 0.6200, 0.7056, 0.2929,  0.6744],                [0.2000, 0.3003, 0.2290, 0.8110, 0.9529, 0.2384,  0.0442],                [0.1000, 0.7809, 0.5374, 0.1381, 0.5655, 0.5679,  0.6123],                [0.7000, 0.3454, 0.9352, 0.0283, 0.7969, 0.7874,  0.0732],                [0.9000, 0.0395, 0.4250, 0.2010, 0.8243, 0.9836, -0.0281],                [0.1000, 0.2542, 0.8055, 0.0806, 0.0381, 0.1833,  0.0791],                [0.2000, 0.8412, 0.2388, 0.0388, 0.8542, 0.4247,  0.1126],                [0.8000, 0.7370, 0.3922, 0.5172, 0.3952, 0.6280,  0.2454],                [0.6000, 0.1313, 0.5940, 0.6993, 0.2744, 0.0337, -0.0489],                [0.9000, 0.8016, 0.2760, 0.9357, 0.4104, 0.7371,  0.0214],                [0.6000, 0.6316, 0.9115, 0.3722, 0.0273, 0.0787,  0.5736],                [0.7000, 0.8821, 0.1944, 0.1973, 0.6821, 0.2330,  0.0460],                [0.4000, 0.6779, 0.8546, 0.8716, 0.7558, 0.6497,  0.0783],                [1.0000, 0.5682, 0.1289, 0.4361, 0.1452, 0.4574,  0.0586],                [0.3000, 0.2194, 0.6718, 0.2631, 0.4686, 0.3043,  0.6239],                [1.0000, 0.3801, 0.7743, 0.7396, 0.8837, 0.8693, -0.0627],                [0.4000, 0.9939, 0.6117, 0.2750, 0.9850, 0.8189,  0.0045],                [0.6000, 0.4286, 0.9720, 0.3607, 0.5942, 0.4977,  0.2505],                [0.9000, 0.1827, 0.6516, 0.1159, 0.9271, 0.3751,  0.0581],                [0.5000, 0.7077, 0.4601, 0.9771, 0.2619, 0.5356, -0.0289],                [0.0000, 0.4767, 0.3555, 0.6558, 0.1989, 0.0162,  0.1569],                [0.3000, 0.4371, 0.7413, 0.9112, 0.0723, 0.9252,  0.9560],                [0.5000, 0.5276, 0.0843, 0.7266, 0.2118, 0.5302,  0.2011],                [0.7000, 0.1468, 0.3188, 0.3087, 0.3526, 0.3344,  0.5923],                [0.0000, 0.9432, 0.5099, 0.5586, 0.6024, 0.6905, -0.0087],                [0.1000, 0.0802, 0.0017, 0.4008, 0.5019, 0.1411,  0.2529],                [0.8000, 0.0221, 0.7221, 0.8440, 0.1240, 0.9540,  0.2324],                [0.5000, 0.2958, 0.4872, 0.5731, 0.4418, 0.7037,  0.7130],                [0.2000, 0.9085, 0.8757, 0.4770, 0.6390, 0.1237,  1.1281],                [0.7000, 0.0000, 0.3683, 0.0712, 0.0515, 0.4062,  0.1484],                [0.3000, 0.7281, 0.9619, 0.0431, 0.0000, 0.1199,  0.0894],                [0.8000, 1.0000, 0.8457, 1.0000, 1.0000, 0.0000, -0.0373],                [0.8000, 0.9280, 0.8906, 0.4856, 0.6641, 0.1195,  0.1296],                [0.8000, 1.0000, 1.0000, 0.5266, 0.6308, 0.0358,  0.3261],                [1.0000, 0.5787, 0.4309, 0.1291, 0.4129, 0.5590,  0.2360],                [0.3000, 0.0000, 1.0000, 0.5009, 0.6804, 0.1289,  0.0289],                [0.4000, 0.9628, 0.9097, 0.4806, 0.4525, 0.0000,  2.6153],                [0.5000, 0.0000, 1.0000, 1.0000, 1.0000, 1.0000, -0.0507],                [1.0000, 1.0000, 0.0000, 1.0000, 0.0000, 1.0000, -0.1304],                [0.3000, 0.0000, 0.0000, 0.0000, 0.0000, 1.0000,  0.1087],                [0.4000, 1.0000, 1.0000, 0.5046, 1.0000, 0.0000,  2.5231],                [0.4000, 1.0000, 1.0000, 0.4928, 0.0000, 1.0000,  0.0980],                [0.4000, 1.0000, 1.0000, 0.4467, 1.0000, 1.0000, -0.0251],                [0.5000, 1.0000, 0.0000, 0.5280, 0.0000, 0.0000,  2.2291],                [0.4000, 0.9658, 0.0000, 0.4804, 0.0000, 0.0898,  2.5967],                [0.4000, 0.8883, 1.0000, 0.5189, 0.0000, 0.0000,  3.1767],                [0.4000, 1.0000, 0.0000, 0.3745, 1.0000, 0.0000,  1.6312],                [0.7000, 0.0000, 1.0000, 0.5100, 1.0000, 1.0000, -0.0249],                [0.4000, 0.8786, 0.0161, 0.5639, 0.0581, 0.0000,  3.1551],                [0.5000, 0.8284, 1.0000, 0.6011, 0.0000, 0.0000,  2.5988],                [0.4000, 0.9136, 1.0000, 0.5669, 0.0000, 0.0802,  3.2133],                [0.6000, 1.0000, 0.0000, 0.7980, 1.0000, 1.0000, -0.0130],                [0.4000, 0.8908, 1.0000, 0.5868, 0.0000, 0.0000,  3.0625],                [0.4000, 0.8291, 1.0000, 0.5621, 0.0000, 0.0862,  2.8820],                [0.0000, 1.0000, 0.0000, 0.0000, 0.0000, 0.0000,  0.0447],                [0.3000, 0.8511, 1.0000, 0.7110, 0.0000, 0.0000,  2.1879],                [0.4000, 0.8976, 0.4139, 0.5711, 1.0000, 0.0425,  2.9899],                [0.5000, 0.9286, 0.0000, 0.6761, 0.0000, 0.1140,  2.2584],                [0.4000, 0.9195, 0.0000, 0.6349, 0.0000, 0.1203,  2.5906],                [0.1000, 0.2126, 1.0000, 0.5955, 0.0000, 1.0000,  0.4850],                [0.4000, 0.8943, 1.0000, 0.5565, 0.0000, 0.0097,  2.9492],                [0.2000, 0.4618, 0.0000, 0.1665, 1.0000, 1.0000,  0.1278],                [0.2000, 1.0000, 1.0000, 0.9451, 0.0000, 0.0000,  0.4609],                [0.5000, 1.0000, 1.0000, 0.5839, 0.0000, 0.3425,  0.7659],                [0.4000, 0.8899, 1.0000, 0.5622, 0.0000, 0.0421,  3.1330],                [0.4000, 0.8026, 0.0000, 0.5356, 0.0000, 0.0000,  2.7828],                [0.2000, 0.0556, 1.0000, 1.0000, 0.0000, 0.6695,  0.1134],                [0.4000, 0.7124, 1.0000, 0.6842, 0.0000, 0.0000,  2.3028]],                dtype=torch.float64) NUBO efficiently explores the parameter space by switching between exploring areas with high uncertainty and high predictions. This means the algorithm does not monotonically converge to a single solution as conventional optimization algorithms would. Thus, the approximate solution to an objective function is the best value found during optimization. In this case study, the approximate solution, i.e., the solution with the highest output (last column in the Python output above), was found at iteration 53, and the inputs and outputs are printed below.
        >>> best_iter = int(torch.argmax(y_train))        >>>        >>> print("Approximate solution")        >>> print("--------------------")        >>> print(f"Evaluation: {best_iter + 1}")        >>> print(f"Inputs: {x_train[best_iter]}")        >>> print(f"Output: {y_train[best_iter]:.4f}")
@@
        Approximate solution        --------------------        Evaluation: 53        Inputs: tensor([0.4000, 0.9136, 1.0000, 0.5669, 0.0000, 0.0802],                       dtype=torch.float64)        Output: 3.2133 We compare the results provided by NUBO with the results from random sampling and using a space-filling design, in this case, Latin hypercube sampling (LHS). The code below generates results for the total budget of 70 evaluations for both sampling methods and plots the results against each other, with the number of evaluations on the x-axis and the accumulative best output for each method on the y-axis. Figure 5 shows that NUBO (green line) provides a better solution than either alternative approach and is very close to the true maximum of 3.32237. NUBO succeeds in accurately approximating the true optimum.
@@
Figure 5: Results of the Bayesian optimization algorithm implemented with NUBO, as defined in this case study, compared to random sampling and Latin hypercube sampling.
    $>$ import matplotlib.pyplot as plt    $>$ import numpy as np    $>$    $>$    $>$ torch.manual_seed(123)    $>$ random = black_box(torch.rand((70, dims)))    $>$ lhs = black_box(gen_inputs(num_points=70,    ...                            num_dims=dims,    ...                            bounds=bounds))    $>$    $>$ plt.plot(range(1, 71), np.maximum.accumulate(random), label="Random")    $>$ plt.plot(range(1, 71), np.maximum.accumulate(lhs), label="LHS")    $>$ plt.plot(range(1, 71), np.maximum.accumulate(y_train), label="NUBO")    $>$ plt.hlines(3.32237, 0, 71, colors="red", linestyles="dashed",    ...            label="Maximum")    $>$ plt.title("Comparison against random designs")    $>$ plt.xlabel("Evaluations")    $>$ plt.ylabel("Output")    $>$ plt.legend(loc=’lower center’, ncol=4, bbox_to_anchor=(0.5, -0.275))    $>$ plt.xlim(0, 71)    $>$ plt.tight_layout()

### v3

@@ 4 Case study
We present the general workflow for optimising an expensive-to-evaluate black-box function with NUBO by providing a detailed case study in which a test function with six input dimensions is optimised. This case study demonstrates how the user can specify the parameter input space, generate initial training data, and define and run the Bayesian optimisation loop.
@@
    >>> import torch
    >>> torch.manual_seed(123)
    >>> torch.set_printoptions(precision = 4, sci_mode = False)
@@
A typical objective function optimised with Bayesian optimisation is expensive to evaluate and thus not feasible to use in a case study that aims to illustrate how NUBO can be applied. Hence, we will use one of the synthetic test functions provided by NUBO as a surrogate expensive-to-evaluate black-box function. We use the six-dimensional Hartmann function that possesses multiple local and one global minimum. Its input space is bounded by the hyper-rectangle $0,1]^{6}$. Observational noise, such as measurement error, is represented by adding a small amount of random Gaussian noise to the function output by setting noise_std = 0.1. minimise is set to False to transform the minimisation into a maximisation problem as required for Bayesian optimisation with NUBO.
    >>> from nubo.test_functions import Hartmann6D
    >>> black_box = Hartmann6D(noise_std = 0.1, minimise = False)
@@
With our objective function specified, we can focus on defining the input space. We know that our objective function has six inputs that are all bounded by $[0,1]$. As introduced in Section 3.2, the bounds are defined as a $2\times d$ torch.tensor, where the first row specifies the lower bounds and the second row specifies the upper bounds. This case study also highlights the mixed parameter optimisation capabilities of NUBO (see Section 3.2) by assuming that the first input is a discrete parameter restricted to 0.2, 0.4, 0.6, and 0.8. We can implement this by specifying a dictionary, where the key is the input dimension and the value is a list of possible values the input can take, that is {0: [0.0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0]}. Note that indexing starts at zero in Python.
    >>> dims = 6
    >>> bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],
    ...                        [1., 1., 1., 1., 1., 1.]])
    >>> discrete = {0: [0.0, 0.1, 0.2, 0.3, 0.4, 0.5,
    ...                 0.6, 0.7, 0.8, 0.9, 1.0]}
@@
The Bayesian optimisation loop requires initial training data. This is important to train the Gaussian process that tries to emulate the objective function. This case study uses the gen_inputs function introduced in Section 3.3 to generate 30 initial data points from a Latin hypercube design. We round the first input dimension to fit the discrete values specified above as Latin hypercube designs return continuous values. These points are evaluated by the objective function to complete our training data pairs consisting of input parameters x_train and observations y_train.
    >>> from nubo.utils import gen_inputs
    >>> x_train = gen_inputs(num_points = dims * 5,
    ...                      num_dims = dims,
    ...                      bounds = bounds)
    >>> x_train[:, 0] = torch.round(x_train[:, 0], decimals = 1)
    >>> y_train = black_box(x_train)
@@
Next, we specify the Bayesian optimisation algorithm we plan to use in our optimisation loop. We define the bo function that takes our training pairs (x_train, y_train) and returns the next candidate point x_new which is evaluated by the objective function in four steps. First, we set up our surrogate model as the Gaussian process provided by NUBO with a Gaussian likelihood as discussed in Section 3.1. Second, we train the Gaussian process gp with our training data by maximising the likelihood with the Adam algorithm (Kingma and Ba, 2014) via the fit_gp function. Here, we set a custom learning rate lr and the number of optimisation steps steps. Third, we define an acquisition function that will guide our optimisation. As we assume that our objective function allows parallel function evaluations, we decide to compute multi-point batches at each iteration and choose a Monte Carlo acquisition function, in this case MCUpperConfidenceBound. The acquisition function acq is instantiated by providing it with the fitted Gaussian process gp, a value for the trade-off hyper-parameter beta, and the number of Monte Carlo samples used to approximate the acquisition function. For further details, refer to Section 2.2. Fourth, we maximise the acquisition function acq with the multi_sequential function that uses the sequential strategy for computing multiple candidate points. We decide to compute four candidate points at each iteration by setting batch_size=4 and providing the previously specified bounds and discrete values. The Adam optimiser is used as Monte Carlo acquisition functions require a stochastic optimiser due to their inherent randomness introduced by drawing the Monte Carlo samples. The optimiser is initialised at two different initial points chosen as the two points with the highest acquisition value out of 100 potential points sampled from a Latin hypercube design. We chose two initialisations to keep the computational overhead within the replication script low. In practice, a higher number of initialisations might be beneficial.
    >>> from nubo.acquisition import MCUpperConfidenceBound
    >>> from nubo.models import GaussianProcess, fit_gp
    >>> from nubo.optimisation import multi_sequential
    >>> from gpytorch.likelihoods import GaussianLikelihood
    >>> def bo(x_train, y_train):
    >>>     likelihood = GaussianLikelihood()
    >>>     gp = GaussianProcess(x_train, y_train, likelihood = likelihood)
    >>>     fit_gp(x_train, y_train, gp = gp, likelihood = likelihood,
    ...            lr = 0.1, steps = 200)
    >>>     acq = MCUpperConfidenceBound(gp = gp, beta = 4, samples = 128)
    >>>     x_new, _ = multi_sequential(func = acq,
    ...                                 method = "Adam",
    ...                                 batch_size = 4,
    ...                                 bounds = bounds,
    ...                                 discrete = discrete,
    ...                                 lr = 0.1,
    ...                                 steps = 200,
    ...                                 num_starts = 2,
    ...                                 num_samples = 100)
    >>>     return x_new
@@
Finally, we specify the entire optimisation loop, that is, a simple for-loop that computes the next batch of candidate points using the defined Bayesian optimisation algorithm bo, evaluates the candidate points by the objective function black_box, and adds the new data pairs (x_new, y_new) to the training data. We let the optimisation loop run for ten iterations and print all evaluations, where the first six columns are the inputs and the final column is the output from the objective function. The first 30 rows give the initial training data generated by the Latin hypercube design, while the last 40 rows were chosen by the Bayesian optimisation algorithm. The results clearly show that NUBO improves upon the initial space-filling design and produces points which are consistent with the bounds and discrete values that specify the parameter input space.
        >>> iters = 10
        >>> for iter in range(iters):
        >>>     x_new = bo(x_train, y_train)
        >>>     y_new = black_box(x_new)
        >>>     x_train = torch.vstack((x_train, x_new))
        >>>     y_train = torch.hstack((y_train, y_new))
        >>> print(torch.hstack([x_train, y_train.reshape(-1, 1)]))
        tensor([[0.2000, 0.6523, 0.1574, 0.7822, 0.3039, 0.8603,  0.1251],
                [0.5000, 0.9127, 0.8746, 0.4787, 0.6523, 0.1249,  2.2907],
                [0.4000, 0.5638, 0.0459, 0.6200, 0.7056, 0.2929,  0.6744],
                [0.2000, 0.3003, 0.2290, 0.8110, 0.9529, 0.2384,  0.0442],
                [0.1000, 0.7809, 0.5374, 0.1381, 0.5655, 0.5679,  0.6123],
                [0.5000, 1.0000, 1.0000, 0.5839, 0.0000, 0.3425,  0.7659],
                [0.4000, 0.8899, 1.0000, 0.5622, 0.0000, 0.0421,  3.1330],
                [0.4000, 0.8026, 0.0000, 0.5356, 0.0000, 0.0000,  2.7828],
                [0.2000, 0.0556, 1.0000, 1.0000, 0.0000, 0.6695,  0.1134],
                [0.4000, 0.7124, 1.0000, 0.6842, 0.0000, 0.0000,  2.3028]],
                dtype=torch.float64)
@@
NUBO explores the parameter space efficiently by switching between exploring areas with high uncertainty and areas with high predictions. This means that the algorithm does not monotonically converge to a single solution as conventional optimisation algorithms would. Thus, the approximate solution to an objective function is the best value found during optimisation. In this case study, the approximate solution, i.e., the solution with the highest output (last column in the Python output above), was found at iteration 53, and the inputs and outputs are printed below.
        >>> best_iter = int(torch.argmax(y_train))
        >>> print("Approximate solution")
        >>> print("--------------------")
        >>> print(f"Evaluation: {best_iter + 1}")
        >>> print(f"Inputs: {x_train[best_iter]}")
        >>> print(f"Output: {y_train[best_iter]:.4f}")
            Approximate solution
        Evaluation: 53
        Inputs: tensor([0.4000, 0.9136, 1.0000, 0.5669, 0.0000, 0.0802],
                       dtype=torch.float64)
        Output: 3.2133
@@
We compare the results provided by NUBO with the results from random sampling and using a space-filling design, in this case, Latin hypercube sampling (LHS). The code below generates results for the full budget of 70 evaluations for both sampling methods and plots the results against each other, with the number of evaluations on the x-axis and the accumulative best output for each method on the y-axis. Figure 5 shows that NUBO (green line) provides a better solution than either alternative approach and is very close to the true maximum of 3.32237. NUBO succeeds in accurately approximating the true optimum.
@@
Figure 5: Results of the Bayesian optimisation algorithm implemented with NUBO as defined in this case study compared to random sampling and Latin hypercube sampling.
    >>> import matplotlib.pyplot as plt
    >>> import numpy as np
    >>> torch.manual_seed(123)
    >>> random = black_box(torch.rand((70, dims)))
    >>> lhs = black_box(gen_inputs(num_points = 70,
    ...                            num_dims = dims,
    ...                            bounds = bounds))
    >>> plt.plot(range(1, 71), np.maximum.accumulate(random), label = "Random")
    >>> plt.plot(range(1, 71), np.maximum.accumulate(lhs), label = "LHS")
    >>> plt.plot(range(1, 71), np.maximum.accumulate(y_train), label = "NUBO")
    >>> plt.hlines(3.32237, 0, 71, colors = "red", linestyles = "dashed",
    ...            label = "Maximum")
    >>> plt.title("Comparison against random designs")
    >>> plt.xlabel("Evaluations")
    >>> plt.ylabel("Output")
    >>> plt.legend(loc = ’lower center’, ncol = 4, bbox_to_anchor = (0.5, -0.275))
    >>> plt.xlim(0, 71)
    >>> plt.tight_layout()